What is the initial velocity of an automobile 5.0m/s² and a displacement of 114m


1
Expert's answer
2021-08-16T17:04:02-0400
"s=vt+0.5at^2\\114=6v+0.5(5)(6)^2\\v=4\frac{m}{s}"